PHOTOCHEMISTRY OF BROMOFORM AND TRIBROMIDES OF OTHER ELEMENTS IN SOLUTION

\underline{ANDREY S. MERESHCHENKO}, KANYKEY E. KARABAEVA, ALEXANDER N. TARNOVSKY, Department of Chemistry and Center for Photochemical Sciences, Bowling Green State University, Bowling Green, Ohio 43403; PATRICK Z. EL-KHOURY, Institute for Surface and Interface Science, University of California Irvine, Irvine, CA 92697; AND SUMAN K. PAL, School of Basic Sciences IIT Mandi, Vallabh Degree College Campus, Mandi 175001, India.

Photochemistry of bromoform in solution was studied by means of ultrafast time-resolved transient absorption spectroscopy. After 255 nm excitation, bromoform dissociates to the CHBr_{2} radical and bromine atom, which recombine to form iso-bromoform CHBr_{2}-Br. In nonpolar solvents, such as methylcyclohexane, this isomer has a lifetime significantly greater than time window (1.2 ns), while in polar solvents, such as acetonitrile and methanol, iso-bromoform relaxes to the parent molecule in about 200 ps. This behavior is consistent with DFT intrinsic reaction coordinate calculations of the ground state potential energy surfaces in these solvents. Also, we showed photochemical formation of isomers with Br-Br bond in tribromides of other elements.